# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.2.1. Stability for $\psi_{i}\equiv 0$
(3.104) $\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{1}(\eta_{1}),\end{cases}\,\text{with the following state variables :}\,\begin{cases}\eta_{2},\\ z_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z_{2}=\eta_{4}-\alpha_{2}(z_{1}),\\ z_{3}=\eta_{1}-\alpha_{3}(z_{2}),\end{cases}$
@@
These convex Lyapunov functions, are well suited to study the equilibrium stability and to quantify how the cyclic interactions and control laws contribute to the system’s energy decay. Indeed,
(3.106) $\dot{\omega}_{1}(\eta_{2})=-\phi_{2}(\eta_{2})\phi_{3}(\eta_{3})$
We design the first fictitious control $\eta_{2}$ so that
(3.107) $\phi_{3}(\eta_{2})=c_{1}\phi_{2}(\eta_{2})\Longrightarrow\eta_{2}=\ln\left(1+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\right)$
we have
(3.108) $\underbrace{\lambda_{3}(e^{z_{1}+\eta_{2}}-1)}_{\phi_{3}(\eta_{3})})=\phi_{3}(z_{1})e^{\eta_{2}}+\phi_{3}(\eta_{2})$
and thus
(3.109) $\dot{\omega}^{1}_{1}(\eta_{2})=-c_{1}\phi^{2}_{2}(\eta_{2})-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})$
Consider the second Lyapunov function for $(z_{1},\eta_{2})$
(3.110) $\omega^{1}_{2}(\eta_{2},z_{1})=\omega^{1}_{1}(\eta_{2})+\lambda_{3}(e^{z_{1}}-1-z_{1})$
and
(3.111) $\dot{\omega}^{1}_{2}(\eta_{2},z_{1})=-c_{1}\phi^{2}_{2}(\eta_{2})-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})+\phi_{3}(z_{1})\left(-\phi_{4}(\eta_{4})+\phi_{3}(\eta_{3})\right)$
The second fictitious control $z_{1}$ is given by
(3.112) $\phi_{4}(z_{1})=c_{2}\phi_{3}(z_{1})$
and
(3.113) $\underbrace{\lambda_{4}(e^{z_{2}+z_{1}}-1)}_{\phi_{4}(\eta_{4})})=\phi_{4}(z_{2})e^{z_{1}}+\phi_{4}(z_{1})$
Then,
(3.114) $\dot{\omega}^{1}_{2}(\eta_{2},z_{1})=-c_{1}\phi^{2}_{2}(\eta_{2})-c_{2}\phi^{2}_{3}(z_{1})+\phi^{2}_{3}(z_{1})-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})-\frac{c_{2}}{\lambda_{4}}\phi^{2}_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})$
$+c_{1}\phi_{2}(\eta_{2})\phi_{3}(z_{1})$
Let us consider the third Lyapunov function given by
(3.115) $\omega^{1}_{3}(\eta_{2},z_{1},z_{2})=\omega^{1}_{2}(\eta_{2},z_{1})+\lambda_{4}(e^{z_{2}}-1-z_{2})$
and the fictitious control $z_{2}$ is given by
(3.116) $\phi_{1}(z_{2})=c_{3}\phi_{4}(z_{2}).$
It follows from (3.49) that
(3.117) $\dot{\omega}^{1}_{3}(\eta_{2},z_{1},z_{2})=-c_{1}\phi^{2}_{2}(\eta_{2})-c_{2,o}\phi^{2}_{3}(z_{1})-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})-\frac{c_{2}}{\lambda_{4}}\phi^{2}_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})$
$+c_{1}\phi_{2}(\eta_{2})\phi_{3}(z_{1})+\phi_{4}(z_{2})\dot{z}_{2}$
We compute $\dot{z}_{2}$ in the form
(3.118) $\dot{z}_{2}=-c_{3}\phi_{4}(z_{2})-\phi_{1}(z_{3})-\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})+c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})-c_{1}\phi_{2}(\eta_{2})-\phi_{3}(z_{1})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})$
thanks to
(3.119) $\phi_{1}(\eta_{1})=\phi_{1}(z_{3}+z_{2})=c_{3}\phi_{4}(z_{2})+\phi_{1}(z_{3})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3}).$
Finally, $\dot{\omega}^{1}_{3}(\eta_{2},z_{1},z_{2})$ is given by
(3.120) $\dot{\omega}^{1}_{3}(\eta_{2},z_{1},z_{2})=-c_{1}\phi^{2}_{2}(\eta_{2})-c_{2}\phi^{2}_{3}(z_{1})-c_{3}\phi^{2}_{4}(z_{2})+\phi^{2}_{3}(z_{1})+\phi^{2}_{4}(z_{2})-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})$
$-\frac{c_{2}}{\lambda_{4}}\phi^{2}_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})+c_{1}\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\phi_{1}(z_{3})\phi_{4}(z_{2})-\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})\phi_{1}(z_{3})+c_{2}\phi_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi^{2}_{4}(z_{2})$
$-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})\phi_{4}(z_{2})$
Let the last Lyapunov control function be defined by
(3.121) $V_{4}(\eta_{2},z_{1},z_{2},z_{3})=\theta\omega^{1}_{3}(\eta_{2},z_{1},z_{2})+\lambda_{1}(e^{z_{3}}-1-z_{3})$
and
(3.122) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=\theta\dot{\omega}^{1}_{3}(\eta_{2},z_{1},z_{2})+\phi_{1}(z_{3})\dot{z}_{3}$
We shown that
(3.123) $\dot{z}_{3}=u^{*}-u-\phi_{2}(\eta_{2})+c_{3}\phi_{4}(z_{2})+\phi_{1}(z_{3})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})-c_{2}\phi_{3}(z_{1})-\phi_{4}(z_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})$
$+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})$
and therefore
(3.124) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})$
$+\theta[-\phi_{2}(\eta_{2})\phi_{3}(z_{1})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\phi_{3}(z_{1})-\frac{c_{2}}{\lambda_{4}}\phi^{2}_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})+c_{1}\phi_{2}(\eta_{2})\phi_{3}(z_{1})+\phi^{2}_{3}(z_{1})+\phi^{2}_{4}(z_{2})$
$+c_{2}\phi_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi^{2}_{4}(z_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})-\phi_{3}(z_{1})\phi_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})\phi_{4}(z_{2})]$
$\phi_{1}(z_{3})[(c_{1}-1)\phi_{2}(\eta_{2})+(c_{3}-1-\theta)\phi_{4}(z_{2})-(c_{2}-1)\phi_{3}(z_{1})+\phi_{1}(z_{3})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})]$
$+\phi_{1}(z_{3})[u^{*}-u-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})]$
Then, with a control of the form
@@
(3.125) $\begin{aligned} u&=u^{*}+c_{40}\phi_{1}(z_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})+(c_{3}-1-\theta)\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})-(c_{2}-1)\phi_{3}(z_{1})\\ &+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})+\frac{\theta}{\phi_{1}(z_{3})}\left(\phi^{2}_{4}(z_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})+\phi^{2}_{3}(z_{1})\right)\\ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{3})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z_{2})\right),\end{aligned}$
we finally obtain
(3.126) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})-c_{4}\phi^{2}_{1}(z_{3}).$
@@ 3.2.2. Stability for $\psi_{i}\neq 0$
(3.128) $V_{G}(\eta,\psi)=V_{4}(\eta_{2},z_{1},z_{2},z_{3})\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\frac{\gamma_{3}}{\sigma_{3}}h(G_{3}(\psi_{3}))\;+\;\frac{\gamma_{4}}{\sigma_{4}}h(G_{4}(\psi_{4})).$
Considering the Lyapunov control function (3.121), and by following the approach of Section 3.1.1 with control (3.125) applied to system (3.127), we obtain
(3.129) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})-c_{4}\phi^{2}_{1}(z_{3})$
$\underbrace{\left(\phi_{1}(z_{3})-\theta\phi_{4}(z_{2})\right)}_{A_{1}}|\hat{\phi}_{1}-\phi_{1}(\eta_{1})|\underbrace{-\phi_{1}(z_{3})}_{A_{2}}|\hat{\phi}_{2}-\phi_{2}(\eta_{2})|+\underbrace{\left(\phi_{1}(z_{3})-\theta\phi_{4}(z_{2})+\theta\phi_{3}(z_{1})-\theta\phi_{2}(\eta_{2})\right)}_{A_{3}}|\hat{\phi}_{3}-\phi_{3}(\eta_{3})|$
$+\underbrace{\left(\theta\phi_{4}(z_{2})-\phi_{1}(z_{3})-\theta\phi_{3}(z_{1})\right)}_{A_{4}}|\hat{\phi}_{4}-\phi_{4}(\eta_{4})|+\mathcal{R}_{1}$
where
(3.130) $\mathcal{R}_{1}=\phi_{1}(z_{3})(-\phi_{1}(z_{3})+\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})-(c_{3}-1-\theta)\phi_{4}(z_{2})-(c_{1}-1)\phi_{2}(\eta_{2})-\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+(c_{2}-1)\phi_{3}(z_{1}))$
$-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1}))+\theta c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})+\theta c_{1}\phi^{2}_{2}(\eta_{2})+\theta c_{2}\phi^{2}_{3}(z_{1})-\theta\phi^{2}_{3}(z_{1})+\theta c_{3}\phi^{2}_{4}(z_{2})-\theta\phi^{2}_{4}(z_{2})$
$-\theta\phi_{3}(z_{1})\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z_{2})\right)$
$+\left(\phi_{1}(z_{3})-\theta\phi_{4}(z_{2})\right)\phi_{1}(\eta_{1})-\phi_{1}(z_{3})\phi_{2}(\eta_{2})+\left(\phi_{1}(z_{3})-\theta\phi_{4}(z_{2})+\theta\phi_{3}(z_{1})-\theta\phi_{2}(\eta_{2})\right)\phi_{3}(\eta_{3})+\left(\theta\phi_{4}(z_{2})-\phi_{1}(z_{3})-\theta\phi_{3}(z_{1})\right)\phi_{4}(\eta_{4}).$
With
(3.131) $\begin{cases}\phi_{1}(\eta_{1})=c_{3}\phi_{4}(z_{2})+\phi_{1}(z_{3})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3}),\\ \phi_{3}(\eta_{3})=c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2}),\\ \phi_{4}(\eta_{4})=c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2}),\end{cases}$
from (3.108)-(3.113)-(3.119), we get
(3.132) $\mathcal{R}_{1}=0.$
Then
(3.133) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})-c_{4}\phi^{2}_{1}(z_{3})$
(3.134) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})-c_{4}\phi^{2}_{1}(z_{3})+\left(C_{1}|\phi_{1}(\eta_{1})+\lambda_{1}|-\gamma_{1}\right)(e^{G_{1}}-1)$
Selecting $\eta$ as follows
(3.135) $\begin{cases}\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\\ \\ \eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\\ \\ \eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\\ \\ \eta_{4}\leq\ln\left(\frac{\gamma_{4}}{C_{4}\lambda_{4}}\right),\end{cases}\qquad\text{with}\;\begin{cases}\gamma_{1}>C_{1}\lambda_{1},\\ \\ \gamma_{2}>C_{2}\lambda_{2},\\ \\ \gamma_{3}>C_{3}\lambda_{3},\\ \\ \gamma_{4}>C_{4}\lambda_{4},\end{cases}$
yields the desired solution.
@@
(3.136) $\mathcal{K}_{4}=\left\{(\eta,\psi)\in\mathbb{R}^{4}\times\mathcal{S}^{4}\ \middle|\ \begin{aligned} &\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\qquad&&\gamma_{1}>C_{1}\lambda_{1},\\[4.0pt] &\eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\qquad&&\gamma_{2}>C_{2}\lambda_{2},\\[4.0pt] &\eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\qquad&&\gamma_{3}>C_{3}\lambda_{3},\\[4.0pt] &\eta_{4}\leq\ln\left(\frac{\gamma_{4}}{C_{4}\lambda_{4}}\right),\qquad&&\gamma_{4}>C_{4}\lambda_{4},\\[6.0pt] &u^{*}+c_{40}\phi_{1}(z_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})+(c_{3}-1-\theta)\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})\\ &+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})-(c_{2}-1)\phi_{3}(z_{1})\\ &+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})+\frac{\theta}{\phi_{1}(z_{3})}\left(\phi^{2}_{4}(z_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})+\phi^{2}_{3}(z_{1})\right)\\ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{3})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\right)\\ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{3})}\left((c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z_{2})\right)>0.\end{aligned}\right\}.$

### v2

@@ 3.2.1. Stability for $\psi_{i}\equiv 0$
(3.104) $\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{1}(\eta_{1}),\end{cases}\,\text{with the following state variables :}\,\begin{cases}\eta_{2},\\ z^{4}_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z^{4}_{2}=\eta_{4}-\alpha_{2}(z^{4}_{1}),\\ z^{4}_{3}=\eta_{1}-\alpha_{3}(z^{4}_{2}),\end{cases}$
@@
These convex Lyapunov functions, are well suited to study the equilibrium stability and to quantify how the cyclic interactions and control laws contribute to the system’s energy decay.
Proposition 3.17.
Under the proposed feedback law the system (3.104) is globally asymptotically stable. The feedback control constructed is uniformly bounded, though not necessarily nonnegative. Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})$ within the set
@@
(3.106) $\mathcal{K}^{\prime}_{4}=\left\{\eta\in\mathbb{R}^{4}\ \middle|\ \begin{aligned} &u^{*}+c_{40}\phi_{1}(z^{4}_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z^{4}_{2})+(c_{3}-1-\theta)\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z^{4}_{2})\phi_{1}(z^{4}_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})\\ &-(c_{2}-1)\phi_{3}(z^{4}_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z^{4}_{1})+\frac{\theta}{\phi_{1}(z^{4}_{3})}\left(\phi^{2}_{4}(z^{4}_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})+\phi^{2}_{3}(z^{4}_{1})\right)\\ &+\frac{\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z^{4}_{3})}(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{4}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z^{4}_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z^{4}_{2})\\ &-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2}))>0\end{aligned}\right\}.$
  • Proof  :
    By considering a Lyapunov function of the form
    (3.107) $V_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})=\theta\lambda_{2}\bigl(e^{\eta_{2}}-1-\eta_{2}\bigr)+\theta\lambda_{3}(e^{z^{4}_{1}}-1-z^{4}_{1})+\theta\lambda_{4}(e^{z^{4}_{2}}-1-z^{4}_{2})+\lambda_{1}(e^{z^{4}_{3}}-1-z^{4}_{3}),$
    and applying the same strategy as in Theorem 3.7, we obtain the control
    (3.108) $\begin{aligned} u=&u^{*}+c_{40}\phi_{1}(z^{4}_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z^{4}_{2})+(c_{3}-1-\theta)\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z^{4}_{2})\phi_{1}(z^{4}_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})\\ &-(c_{2}-1)\phi_{3}(z^{4}_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z^{4}_{1})+\frac{\theta}{\phi_{1}(z^{4}_{3})}\left(\phi^{2}_{4}(z^{4}_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})+\phi^{2}_{3}(z_{1})\right)\\ &+\frac{\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z^{4}_{3})}(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{4}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z^{4}_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z^{4}_{2})\\ &-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})),\end{aligned}$
    and the Lyapunov derivative is given by
    (3.109) $\dot{V}_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z^{4}_{1})-\theta c_{3}\phi^{2}_{4}(z^{4}_{2})-c_{4}\phi^{2}_{1}(z^{4}_{3}).$
    For further details, see the Appendix A.
Remark 3.18.
The control in (3.108) may be rewritten as
(3.110) $\begin{aligned} u=&u^{*}+\frac{1}{\phi_{1}(z^{4}_{3})}(c_{4}\phi^{2}_{1}(z^{4}_{3})+\theta c_{1}\phi^{2}_{2}(\eta_{2})+\theta c_{3}\phi^{2}_{4}(z^{4}_{2})+\theta c_{2}\phi^{2}_{3}(z^{4}_{1}))\\ &+\frac{\phi_{1}(z^{4}_{3})-\theta\phi_{4}(z^{4}_{2})}{\phi_{1}(z^{4}_{3})}\phi_{1}(\eta_{1})-\phi_{2}(\eta_{2})+\frac{\phi_{1}(z^{4}_{3})-\theta\phi_{4}(z^{4}_{2})+\theta\phi_{3}(z^{4}_{1})-\theta\phi_{2}(\eta_{2})}{\phi_{1}(z^{4}_{3})}\phi_{3}(\eta_{3})+\frac{\theta\phi_{4}(z^{4}_{2})-\phi_{1}(z^{4}_{3})-\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z^{4}_{3})}\phi_{4}(\eta_{4})\end{aligned}$
@@ 3.2.2. Stability for $\psi_{i}\neq 0$
(3.112) $V_{G}(\eta,\psi)=V_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\frac{\gamma_{3}}{\sigma_{3}}h(G_{3}(\psi_{3}))\;+\;\frac{\gamma_{4}}{\sigma_{4}}h(G_{4}(\psi_{4})).$
@@
(3.113) $\mathcal{K}_{4}=\left\{(\eta,\psi)\in\mathbb{R}^{4}\times\mathcal{S}^{4}\ \middle|\ \begin{aligned} &\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\qquad&&\gamma_{1}>C_{1}\lambda_{1},\\[4.0pt] &\eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\qquad&&\gamma_{2}>C_{2}\lambda_{2},\\[4.0pt] &\eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\qquad&&\gamma_{3}>C_{3}\lambda_{3},\\[4.0pt] &\eta_{4}\leq\ln\left(\frac{\gamma_{4}}{C_{4}\lambda_{4}}\right),\qquad&&\gamma_{4}>C_{4}\lambda_{4},\\[6.0pt] &u^{*}+c_{40}\phi_{1}(z^{4}_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z^{4}_{2})+(c_{3}-1-\theta)\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})\\ &+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z^{4}_{2})\phi_{1}(z^{4}_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})-(c_{2}-1)\phi_{3}(z^{4}_{1})\\ &+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z^{4}_{1})+\frac{\theta}{\phi_{1}(z^{4}_{3})}\left(\phi^{2}_{4}(z^{4}_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})+\phi^{2}_{3}(z^{4}_{1})\right)\\ &+\frac{\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z^{4}_{3})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{4}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})\right)\\ &+\frac{\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z^{4}_{3})}\left((c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z^{4}_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z^{4}_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})\right)>0.\end{aligned}\right\}.$
  • Proof  :
    By applying the same strategy as in Theorem 3.13, we derive the result with the Lyapunov function given in (3.112) and the control in (3.108). See Appendix A for full details.